A (Ni+2) solution is electrolyzed using a current of 1.25A. What mass of (Ni0) plates out in 30 minutes.
1
Expert's answer
2021-09-15T02:42:38-0400

The decay formula can be written as:

"N=N_0\times 2^{-\frac{t}{t_{1/2}}}"

Therefore:

"t=-log_2(\frac{N}{N_0})\times t_{1/2}=-log_2(\frac{0.5g}{100g})\times 20\,minutes\approx153\,minutes"


Answer153 minutes